Find two consecutive even integers such that the square of the larger is 44 greater that the square of the smaller
Answer by ewatrrr(24785)


Hi,
Let x represent the smaller of the two consecutive even integers,
then (x+2) would be the larger
Write as we Read***
(x+2)^2 = x^2 + 44
solving for x
x^2 + 4x + 4 = x^2 + 44
4x = 40
x = 10 10 and 12 are the two consecutive even integers
checking our answer
144 = 100 + 44